# Nearsightedness in Materials with Indirect Band Gap

### v1

@@ 3.1. Nearsightedness
Theorem 3.4 (Local Perturbations). Suppose $u\in\mathrm{Adm}(\Lambda)$ and ${\mathcal{H}}(u)$ satisfies (TB)  with $\varepsilon_{\mathrm{F}}\notin\sigma({\mathcal{H}}(u))$. Then, there exist positive constants $C_{1},C_{\ell k}(u),\eta_{+}$ and $\eta_{-}$ such that
$\big{|}\rho(u)_{\ell k,ab}\big{|}\leq{\frac{C_{1}}{\mathsf{gap}_{+}}}e^{-\eta_{+}r_{\ell k}}+C_{\ell k}(u)e^{-\eta_{-}r_{\ell k}},$ (3.6)
for all $\ell,k\in\Lambda$ and $1\leq a,b\leq N_{\mathrm{b}}$. Here, $C_{1},\eta_{+}$ are defined in Section 3.1, and ${\eta_{-}}:=c_{0}\min\{h_{0},\gamma_{0}^{d}\mathsf{gap}_{-}\}$ for some $c_{0}>0$ depending only on $d,\mathfrak{m},h_{0}$ and $\gamma_{0}$. Moreover, for sufficiently large $R^{\prime}>0$,
$C_{\ell k}(u):={C(u)}\big{(}\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R^{\prime}})}+e^{-2\eta_{-}\min_{m\in B_{R^{\prime}}}(r_{\ell m}+r_{mk})}\big{)}$ (3.7)
where $C(u):=\frac{C_{2}\|Du\|_{\ell^{2}_{\Upsilon}}}{\mathsf{gap}_{-}}\big{(}1+\frac{R^{d}e^{4\eta_{-}R}}{\mathsf{gap}_{-}^{2d+4}}\big{)}$ with $C_{2}>0$ depending on $h_{0},\gamma_{0},\mathfrak{m},d,N_{\mathrm{b}},M$, and $R>0$ satisfying $\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\setminus B_{R})}\leq C_{3}\mathsf{gap}_{-}$ for some $C_{3}>0$. In particular, for sufficiently small $\|Du\|_{\ell^{2}_{\Upsilon}}$, $R$ and $R^{\prime}$ can be chosen arbitrarily.
Remark 3.5. If $|Du(\ell)|_{\Upsilon}\lesssim(1+|\ell|)^{-\alpha}$ for some $\alpha>d$, we can take $\mathscr{C}$ in (3.8) satisfying $\mathrm{dist}(\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}}))\geq\frac{1}{2}\mathsf{gap}_{-}$, and then obtain a simpler prefactor $C(u)$ of the form depending on $\|Du\|_{\ell^{2}_{\Upsilon}}$ and $\mathsf{gap}_{-}$. ∎
@@
We separate $\left|\rho(u)_{\ell k,ab}\right|$ into two parts by the triangle inequality,
$\left|\rho(u)_{\ell k,ab}\right|\leq\left|\rho^{\mathrm{ref}}_{\ell k,ab}\right|+\left|\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\right|.$
Notice that the first term on the right hand side has already been provided by Section 3.1, so we only need to consider the second term, which can be written as
$\left|\rho(u)_{\ell k,ab}-\rho^{\mathrm{ref}}_{\ell k,ab}\right|=\left|\oint_{\mathscr{C}}\left[\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\big{(}{\mathcal{H}}(u)-{\mathcal{H}}^{\mathrm{ref}}\big{)}\big{(}z-{\mathcal{H}}^{\mathrm{ref}}\big{)}^{-1}\right]_{\ell k,ab}\frac{{\,{\rm d}}z}{2\pi i}\right|.$ (3.8)
For the reference resolvent, we apply the Combes–Thomas estimate (3.4). By approximating ${\mathcal{H}}(u)$ with a finite rank update of ${\mathcal{H}}^{\mathrm{ref}}$, we are able to extend (3.4) as:
$\Big{|}\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}_{\ell k,ab}\Big{|}\leq\frac{4}{\mathfrak{d}^{\mathrm{ref}}}e^{-\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})r_{\ell k}}+C(u)e^{-\frac{1}{2}\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})(|\ell|+|k|)},$
where $\mathfrak{d}^{\rm ref}\coloneqq\mathrm{dist}\big{(}z,\sigma({\mathcal{H}}^{\mathrm{ref}})\big{)}>0$, $\gamma_{\rm{CT}}(\frac{1}{2}\mathfrak{d}^{\mathrm{ref}})$ is defined as in Remark 3.1, and $C(u)$ depends on $\|Du\|_{\ell^{2}_{\Upsilon}}$ and $\mathfrak{d}(u)$. By choosing $\mathscr{C}$ such that $\mathrm{dist}\left(\mathscr{C},\sigma({\mathcal{H}}^{\mathrm{ref}})\right)\geq\frac{1}{2}\mathsf{gap}_{-}$, we can achieve the constant $\eta_{-}$ in the above estimate. For full details, see Section 6.2. ∎
Remark 3.6. The perturbation introduces a site-dependent constant $C_{\ell k}(u)$, which can be arbitrarily small as $|\ell|,|k|\to\infty$. That is, as $\ell,k$ move far away from the perturbation, the estimate in (3.6) will be independent of the lattice sites. Moreover, as $\|Du\|_{\ell^{2}_{\Upsilon}}\to 0$, the estimate converges to the homogeneous case in Section 3.1. Here we also note that the exponent $\eta_{-}$ will be determined by $\mathsf{gap}_{-}$ when the system has a relatively small band gap. ∎
@@
is sufficiently small  [6, 27]. The additional term in (3.6) can be replaced by a site-independent estimate $C(u)e^{-\eta_{-}r_{\ell k}}$ for some constant $C(u)>0$ depending on $\|Du\|_{\ell_{\infty}}$ and $\mathfrak{d}(u)$. ∎
Remark 3.8. More generally, we suppose $O:\mathbb{R}\to\mathbb{R}$ is a function that extends analytically to an open neighbourhood of the contour $\mathscr{C}$. Then, one can define the corresponding local quantities for all $\ell\in\Lambda$ by
$O_{\ell}(u):=\sum_{a=1}^{N_{\mathrm{b}}}\oint_{\mathscr{C}}O(z)\big{(}z-{\mathcal{H}}(u)\big{)}_{\ell\ell,aa}^{-1}\frac{\mathrm{d}z}{2\pi i}.$ (3.9)
Then, the observable corresponding to $O$ is given by $\sum_{\ell}O_{\ell}(u):=\mathrm{Tr}\left(O\big{(}{\mathcal{H}}(u)\big{)}\right)$. All of our results can be extended to a general case in the similar manner, which also improve the results of [5, 27]. ∎
3.2. Strong Locality
The nearsightedness in density matrix is insufficient when we aim to understand locality of interatomic interaction, e.g. to study the construction of interatomic potential models or multi-scale schemes such as QM/MM coupling methods [3, 5, 6, 7, 11], which involve choosing cut-off radius for the interaction. In such cases, the strong locality is required for justification.
The goal of this section is to determine the dependence of strong locality, by providing a sharp estimate of the derivative of the density matrix. In what follows, we will discuss in homogeneous and perturbed systems as before. Since the essence of both is not significantly different, we present the combined results in the following proposition. The outcome of this analysis is that $\mathsf{gap}_{+}\gg\mathsf{gap}_{-}$ does not appear to qualitatively improve the strong locality over previous results. The proof does not give new insights over previous results such as [5, 33] and is therefore postponed to Section 6.3. Instead, we provide a heuristic argument in Remark 3.2 why no better result can be expected. Our numerical experiments provide additional empirical support.
Proposition 3.9 (Locality of the Derivative). Suppose $u\in\mathrm{Adm}(\Lambda)$ and ${\mathcal{H}}(u)$ satisfies (TB) with $\varepsilon_{\mathrm{F}}\notin\sigma({\mathcal{H}}(u))$. Then, for $1\leq i_{1}\leq d$, there exist positive constants $C_{q}(\mathsf{gap}_{-}),C(u)$ and $\eta_{-}$ such that
$\left|\frac{\partial[\rho(u)]_{\ell\ell,ab}}{\partial[u(k)]_{i_{1}}}\right|\leq\left[\frac{C_{q}(\mathsf{gap}_{-})}{(1+r_{\ell k})^{q}}+C(u)\right]^{2}e^{-2\eta_{-}r_{\ell k}},$ (3.10)
for all $\ell\neq k,\ell,k\in\Lambda$ and $1\leq a,b\leq N_{\mathrm{b}}$. Here, $\eta_{-}$ is defined in Section 3.1, and $C_{q}(\mathsf{gap}_{-}):=cq!\frac{\mathsf{gap}_{-}^{q}-1}{\mathsf{gap}_{-}^{q+1}(\mathsf{gap}_{-}-1)}$ for some constant $c>0$ only depending on $d,\mathfrak{m},h_{0},\gamma_{0},N_{\mathrm{b}},M$, and $C(u)$ depends on $\|Du\|_{\ell^{2}_{\Upsilon}}$ and $\mathfrak{d}(u)$.
Remark 3.10. Here, we argue that one may not improve (3.10) by replacing the exponent with a constant depending instead on $\mathsf{gap}_{+}$. First, we note that
$\frac{\partial\rho_{\ell\ell,ab}}{\partial[\bm{r}_{k}]_{j}}$ $\coloneqq\frac{\partial\big{[}\rho(u)\big{]}_{\ell\ell,ab}}{\partial[u(k)]_{j}}\Bigg{|}_{u=0}$
@@ 3.2. Strong Locality
$=\oint_{\mathscr{C}}\left[\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\,\frac{\partial{\mathcal{H}}(u)}{\partial[u(k)]_{i}}\,\big{(}z-{\mathcal{H}}(u)\big{)}^{-1}\right]_{\ell\ell,ab}\frac{{\,{\rm d}}z}{2\pi i}$
@@
$=\sum_{mn,a^{\prime}b^{\prime}}\frac{\partial{\mathcal{H}}_{mn}^{a^{\prime}b^{\prime}}}{\partial[\bm{r}_{k}]_{i}}\fint_{(\Gamma^{\star})^{2}}\left[\oint_{\mathscr{C}}\big{(}z-{\mathcal{H}}_{\xi}\big{)}^{-1}_{\ell_{0}m_{0},aa^{\prime}}\big{(}z-{\mathcal{H}}_{\xi}\big{)}^{-1}_{n_{0}\ell_{0},b^{\prime}b}\frac{{\,{\rm d}}z}{2\pi i}\right]e^{i\xi\cdot(\ell-m)}e^{i\zeta\cdot(n-\ell)}{\,{\rm d}}\xi{\,{\rm d}}\zeta.$ (3.11)
@@
Therefore, as in the proof of Theorem 3.1, one may obtain the strong locality estimate from the locality of the Hamiltonian, as in (TB), and the analyticity of the functions
$(\xi,\zeta)\mapsto\oint_{\mathscr{C}}\big{(}z-{\mathcal{H}}_{\xi}\big{)}^{-1}_{\ell_{0}m_{0},aa^{\prime}}\big{(}z-{\mathcal{H}}_{\zeta}\big{)}^{-1}_{n_{0}\ell_{0},b^{\prime}b}\frac{{\,{\rm d}}z}{2\pi i}.$ (3.12)
@@
Recall that in the analogous function $\widehat{\rho}(\bm{\xi})$ (3.1) for the nearsightedness estimate, one may extend the region of analyticity of $\widehat{\rho}$ to a strip of width proportional to $\mathsf{gap}_{+}$ by replacing $\mathscr{C}$ with a contour depending on $\xi$.
@@
However, this is not the case for the functions (3.12). Since the contour $\mathscr{C}=\mathscr{C}(\xi,\zeta)$ must separate $\bigcup_{n\leq N_{0}}\big{\{}\varepsilon_{n}(\xi)\big{\}}$ from $\bigcup_{n\geq N_{0}+1}\big{\{}\varepsilon_{n}(\zeta)\big{\}}$ and separate $\bigcup_{n\leq N_{0}}\big{\{}\varepsilon_{n}(\zeta)\big{\}}$ from $\bigcup_{n\geq N_{0}+1}\big{\{}\varepsilon_{n}(\xi)\big{\}}$, the region of analyticity of (3.12) is a subset of
@@
$\left\{(\xi_{1},\xi_{2})\in(\mathbb{C}^{d})^{2}\colon\varepsilon_{n}(\xi_{l})\not=\varepsilon_{m}(\xi_{l^{\prime}})\text{ for all }n\leq N_{0}<m\text{ and }l,l^{\prime}=1,2\right\}.$ (3.13)
This region depends on $\mathsf{gap}_{-}$. ∎
@@
In §4, we give numerical experiments in which strongly suggest that the strong locality does indeed depend on $\mathsf{gap}_{-}$ rather than $\mathsf{gap}_{+}$.

### v2

@@ 3. Main Results
Theorem 3.2 (Local perturbations). Suppose that $u\in\mathrm{Adm}(\Lambda)$ and that ${\mathcal{H}}(u)$ satisfies assumption (TB)  with $\varepsilon_{\mathrm{F}}\notin\sigma({\mathcal{H}}(u))$. Then, there exist constants $C_{1},C_{\ell k}(u),\eta_{+},\eta_{-}>0$ such that
(3.6) $\big|\rho(u)_{\ell k}\big|\leq{\frac{C_{1}}{\mathsf{gap}_{+}}}e^{-\eta_{+}r_{\ell k}}+C_{\ell k}(u)e^{-\eta_{-}r_{\ell k}},\quad\forall\ell,k\in\Lambda,$
where $\eta_{+}:=c_{1}\frac{\gamma_{0}}{h_{0}}\min\left\{h_{0},\gamma_{0}^{d}\,\mathsf{gap}_{+}\right\}$ and $\eta_{-}\coloneqq c_{2}\frac{\gamma_{0}}{h_{0}}\min\{h_{0},\gamma_{0}^{d}\mathsf{gap}_{-}\}$. Here, the constants $c_{1},c_{2},C_{1}>0$ depend only on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$, and $C_{\ell k}(u)$ depends on the norm $\|Du\|_{\ell^{2}_{\Upsilon}}$ and satisfies $C_{\ell k}(u)\to 0$ as $|\ell|+|k|-r_{\ell k}\to\infty$.
Remark 3.3. The perturbation induces a site-dependent constant $C_{\ell k}(u)$, which satisfies
$C_{\ell k}(u)\lesssim\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda\backslash B_{R})}+\|Du\|_{\ell^{2}_{\Upsilon}(B_{R})}e^{\eta_{-}R}e^{-\eta_{-}(|\ell|+|k|-r_{\ell k})}$
for some sufficiently large $R>0$. Consequently, in the asymptotic regime where $|\ell|+|k|-r_{\ell k}\to\infty$, the constant $C_{\ell k}(u)$ becomes asymptotically independent of the lattice sites. Moreover, in the limit of vanishing perturbation (i.e., $\|Du\|_{\ell^{2}_{\Upsilon}(\Lambda)}\to 0$), the above estimate converges to the homogeneous bound stated in Section 3. ∎
@@
We split $\left|\rho(u)_{\ell k}\right|$ into two parts by the triangle inequality
$\left|\rho(u)_{\ell k}\right|\leq\left|\rho^{\mathrm{ref}}_{\ell k}\right|+\left|\rho(u)_{\ell k}-\rho^{\mathrm{ref}}_{\ell k}\right|,$
where we have
(3.7) $\left|\rho(u)_{\ell k}-\rho^{\mathrm{ref}}_{\ell k}\right|=\left|\oint_{\mathscr{C}}\left[\big(z-{\mathcal{H}}(u)\big)^{-1}\big({\mathcal{H}}(u)-{\mathcal{H}}^{\mathrm{ref}}\big)\big(z-{\mathcal{H}}^{\mathrm{ref}}\big)^{-1}\right]_{\ell k}\frac{{\,{\rm d}}z}{2\pi i}\right|.$
Then apply the Combes–Thomas estimate (3.4) and its generalization in (3.7) to conclude. For full details, see Section 6.2. ∎
@@
is sufficiently small  [7, 28]. The additional term in (3.6) can be replaced by a site-independent estimate $C(u)e^{-\eta_{-}r_{\ell k}}$ for some constant $C(u)>0$ depending on $\|Du\|_{\ell_{\infty}}$. ∎
As noted previously, there are two types of locality: weak and strong. While the nearsightedness reflecting weak locality of the density matrix is well understood, it is generally insufficient for characterizing the locality of interatomic interactions. This is important in the construction of interatomic potential models or multi-scale schemes such as QM/MM coupling methods [4, 7, 8, 6, 12], where selecting an appropriate interaction cut-off radius is essential, for example. To justify those schemes, the strong locality condition is required.
Following  [7, 28], one can derive a $\mathsf{gap}_{-}$ dependent estimate for the derivative of the density matrix, that is, for $u\in\mathrm{Adm}(\Lambda)$, $1\leq j\leq d$, we have
(3.8) $\left|\frac{\partial[\rho(u)]_{\ell\ell}}{\partial[u(k)]_{j}}\bigg|_{u=\mathbf{0}}\right|\lesssim e^{-\eta_{-}r_{\ell k}},\quad\forall\ell,k\in\Lambda,$
where $\eta_{-}$ is introduced in (3.6).
A natural question is whether the strong locality can also be improved with respect to $\mathsf{gap}_{+}$. Unfortunately, the answer is negative. Here, we provide a heuristic argument explaining why no stronger result can be expected. First, we note that
@@
$\frac{\partial\rho_{\ell\ell}}{\partial[\bm{r}_{k}]_{j}}$ $\coloneqq\frac{\partial\big[\rho(u)\big]_{\ell\ell}}{\partial[u(k)]_{j}}\Bigg|_{u=\mathbf{0}}=\oint_{\mathscr{C}}\left[\big(z-{\mathcal{H}}^{\rm ref}\big)^{-1}\,\frac{\partial{\mathcal{H}}(u)}{\partial[u(k)]_{j}}\bigg|_{u=\mathbf{0}}\,\big(z-{\mathcal{H}}^{\rm ref}\big)^{-1}\right]_{\ell\ell}\frac{{\,{\rm d}}z}{2\pi i}$
@@
(3.9) $=\sum_{mn}\frac{\partial{\mathcal{H}}_{mn}}{\partial[\bm{r}_{k}]_{j}}:\fint_{(\Gamma^{*})^{2}}\left[\oint_{\mathscr{C}}\big(z-{\mathcal{H}}_{\bm{\xi}}\big)^{-1}_{\ell_{0}m_{0}}\big(z-{\mathcal{H}}_{\bm{\zeta}}\big)^{-1}_{n_{0}\ell_{0}}\frac{{\,{\rm d}}z}{2\pi i}\right]e^{i\bm{\xi}\cdot(\ell-m)}e^{i\bm{\zeta}\cdot(n-\ell)}{\,{\rm d}}\bm{\xi}{\,{\rm d}}\bm{\zeta}.$
@@
Therefore, as in the proof of Theorem 3, one may obtain the strong locality estimate from the locality of the Hamiltonian, as in (TB), and the analyticity of the functions
(3.10) $(\bm{\xi},\bm{\zeta})\mapsto\oint_{\mathscr{C}}\big(z-{\mathcal{H}}_{\bm{\xi}}\big)^{-1}_{\ell_{0}m_{0}}\big(z-{\mathcal{H}}_{\bm{\zeta}}\big)^{-1}_{n_{0}\ell_{0}}\frac{{\,{\rm d}}z}{2\pi i}.$
@@
Recall that, for the analogous function $\widehat{\rho}(\bm{\xi})$ (3) used in the nearsightedness estimate, one may extend the region of analyticity of $\widehat{\rho}$ to a strip of width proportional to $\mathsf{gap}_{+}$ by replacing $\mathscr{C}$ with a contour depending on $\bm{\xi}$.
@@
However, this is not the case for the functions (3.10). Since the contour $\mathscr{C}=\mathscr{C}(\bm{\xi},\bm{\zeta})$ must simultaneously separate $\bigcup_{n\leq N_{0}}\big\{\varepsilon_{n}(\bm{\xi})\big\}$ from $\bigcup_{n\geq N_{0}+1}\big\{\varepsilon_{n}(\bm{\xi})\big\}$ and $\bigcup_{n\leq N_{0}}\big\{\varepsilon_{n}(\bm{\zeta})\big\}$ from $\bigcup_{n\geq N_{0}+1}\big\{\varepsilon_{n}(\bm{\zeta})\big\}$, the region of analyticity of (3.10) becomes a subset of
@@
(3.11) $\left\{({\bm{\eta}}_{1},{\bm{\eta}}_{2})\in(\mathbb{R}^{d})^{2}\colon\varepsilon_{N_{0}}(\bm{\xi}+i{\bm{\eta}}_{1})\not=\varepsilon_{N_{0}+1}(\bm{\zeta}+i{\bm{\eta}}_{2}),~\forall(\bm{\xi},\bm{\zeta})\in(\Gamma^{*})^{2}\right\},$
@@
and this region depends on $\mathsf{gap}_{-}$. Additional empirical evidence from our numerical experiments (see Figure 10 and Figure 12 in Section 4), also strongly supports that the strong locality does indeed depend on $\mathsf{gap}_{-}$ rather than $\mathsf{gap}_{+}$.
Nevertheless, we show that the dependence of the strong locality estimate on the spectral gap can be further improved in the small-gap regime, provided that the energy bands near the Fermi level satisfy additional regularity conditions. More precisely, under suitable analyticity assumptions, the decay rate improves from a linear dependence on $\mathsf{gap}_{-}$ to a square-root scaling $\sqrt{\mathsf{gap}_{-}}$.
Specifically, we impose the following regularity assumption on the relevant energy bands:
(Reg) Suppose that the bands $\varepsilon_{N_{0}}$ and $\varepsilon_{N_{0}+1}$ admit an analytic continuation to the domain
$\mathbb{C}_{\delta_{*}}^{d}\coloneqq\{\bm{\xi}+i\bm{\zeta}\in\mathbb{C}^{d}:|\bm{\zeta}|<\delta_{*}\}$
for some constant $\delta_{*}>0$.
Under this assumption, we obtain the following improved bound:
Theorem 3.3. Suppose that $u\in\mathrm{Adm}(\Lambda)$ and that ${\mathcal{H}}^{\rm ref},{\mathcal{H}}(u)$ satisfies assumptions (TB) and (Reg), with $\varepsilon_{\mathrm{F}}\notin\sigma({\mathcal{H}}(u))$. Define
$m_{\infty}\coloneqq\max_{n\in\{N_{0},N_{0}+1\}}\sup_{{\bm{\omega}}\in\mathbb{C}_{\delta_{*}/2}^{d}\cap\mathbb{C}_{\gamma_{0}/2}^{d}}\|\nabla^{2}\varepsilon_{n}({\bm{\omega}})\|_{\rm F}.$
Then,
  • (i)
    (Nearsightedness) There exist constants $C_{2},\widetilde{\eta}_{+}>0$ such that
    (3.12) $\left|\rho^{\mathrm{ref}}_{\ell k}\right|\leq\frac{C_{2}}{\mathsf{gap}_{+}}e^{-\widetilde{\eta}_{+}r_{\ell k}},\quad\forall\ell,k\in\Lambda;$
  • (ii)
    (Strong locality) Moreover, there exist constants $C_{3},\widetilde{\eta}_{-}>0$ such that, for any $1\leq j\leq d$,
    (3.13) $\left|\frac{\partial[\rho(u)]_{\ell\ell}}{\partial[u(k)]_{j}}{\bigg|_{u={\bf 0}}}\right|\leq\frac{C_{3}}{\mathsf{gap}_{-}^{2}}e^{-\widetilde{\eta}_{-}r_{\ell k}},\quad\forall\ell,k\in\Lambda,$
where $\widetilde{\eta}_{+}\coloneqq c_{3}\min\left\{\gamma_{0},\delta_{*},\,\sqrt{\mathsf{gap}_{+}/m_{\infty}}\right\}$ and $\widetilde{\eta}_{-}\coloneqq c_{4}\min\left\{\gamma_{0},\delta_{*},\,\sqrt{\mathsf{gap}_{-}/m_{\infty}}\right\}$ for constants $c_{3},c_{4}>0$. Here, $C_{2}>0$ depends only on $d,\mathfrak{m},M$ and $N_{\mathrm{b}}$, and $C_{3}>0$ depends on $h_{0},\gamma_{0},d,\mathfrak{m},M$ and $N_{\mathrm{b}}$.
Sketch of the proof.
To characterize a suitable analytic subregion of (3.11), it suffices to analyze the behavior of the real parts of the relevant energy bands. By the analyticity assumption (Reg), the band functions $\varepsilon_{n}$ admit a uniform bound on their Hessians over $\mathbb{C}_{\delta_{*}/2}^{d}$, which ensures that $m_{\infty}<\infty$. Consequently, for any $\bm{\xi}+i\bm{\zeta}\in\mathbb{C}_{\delta_{*}/2}^{d}$ and $n\in\{N_{0},N_{0}+1\}$, we obtain
(3.14) $\operatorname{Re}\varepsilon_{n}(\bm{\xi}+i\bm{\zeta})\in B_{m_{\infty}|\bm{\zeta}|^{2}}\!\left(\varepsilon_{n}(\bm{\xi})\right),$
where the linear term $i\nabla\varepsilon_{n}(\bm{\xi})\cdot\bm{\zeta}$ is purely imaginary and therefore does not contribute to the real parts.
Combining this quadratic control of the band separation with the Paley–Wiener estimate, we conclude that the width of the analytic region for (3.10) scales like $\sqrt{\mathsf{gap}_{-}/m_{\infty}}$ in the small-gap regime, and the desired exponential decay estimate follows. The same argument applies to $\rho^{\rm ref}_{\ell k}$ yielding an analogous exponential decay rate scaling with $\sqrt{\mathsf{gap}_{+}/m_{\infty}}$. For full details, see Section 6.3. ∎
Remark 3.5. The assumption (Reg) is generally quite strong but they do apply to certain cases, such as those shown in Figure 4(a), (b) and Figure 5. This condition may fail when band crossings occur in $d\geq 2$ [26, 30]. A typical case are Dirac cones:
${\mathcal{H}}_{\bm{\xi}}=\begin{pmatrix}0&\xi_{1}-i\xi_{2}\\ \xi_{1}+i\xi_{2}&0\end{pmatrix},\quad\bm{\xi}=(\xi_{1},\xi_{2})\in\Gamma^{*},$
where the energy bands $\varepsilon_{\pm}(\bm{\xi})=\pm\sqrt{\xi_{1}^{2}+\xi_{2}^{2}}$ are not even differentiable at $\bm{\xi}={\bf 0}$. The analysis of more complex systems will be left for our future work. ∎
Remark 3.6 (General observables). More generally, we suppose $O:\mathbb{R}\to\mathbb{R}$ is a function that extends analytically to an open neighbourhood of the spectrum. Then, one can define the corresponding local quantities for all $\ell,k\in\Lambda$ by
$O(u)_{\ell k}:=\oint_{\mathscr{C}}O(z)\big(z-{\mathcal{H}}(u)\big)_{\ell k}^{-1}\frac{\mathrm{d}z}{2\pi i},$
and the observable corresponding to $O$ is given by $\sum_{\ell}O(u)_{\ell\ell}:=\mathrm{Tr}\left(O\big({\mathcal{H}}(u)\big)\right)$. The derivatives of $O_{\ell\ell}$ related to mechanical response are given by
$\frac{\partial O(u)_{\ell\ell}}{\partial[u(k)]_{j}}\bigg|_{u={\bf 0}}=\oint_{\mathscr{C}}O(z)\left[(z-{\mathcal{H}}^{\rm ref})^{-1}\frac{\partial{\mathcal{H}}(u)}{\partial[u(k)]_{j}}\bigg|_{u={\bf 0}}(z-{\mathcal{H}}^{\rm ref})^{-1}\right]_{\ell\ell}\frac{{\,{\rm d}}z}{2\pi i}.$
All of our conclusions can be extended to this more general case, which also improve the results of [7, 28]. ∎